The powdered iron reacts spectacularly with chlorine gas, producing a lot of heat, flame, and an iron compound (III).
4Fe - 3Cl2 -- 2Fe2Cl3
a) If 5.00 grams of iron and 10.00 grams of chlorine are available, show which is the limiting reagent
b) Calculate the maximum mass of the product that can be made.
1
Expert's answer
2020-06-09T07:24:29-0400

The equation written here is incorrect as the compound Fe2Cl3 does not exist. The formula of iron (III) chloride is FeCl3, and the correct equation is:

2Fe + 3Cl2 --> 2FeCl3.

n(Fe) = 5.00g / 55.85g/mol = 0.09 mol

n(Cl2) = 10.00g / 70.90g/mol = 0.14 mol

According to the equation, the Fe : Cl2 mole ratio is 2 : 3.

The actual ratio is 2 : ((0.14 x 2) / 0.09) = 2 : 3.1. Therefore Fe is a limiting reactant.

n(FeCl3) = n(Fe) = 0.09 mol

m(FeCl3) = 0.09mol x 162.2g/mol = 14.60 g

Answer: limiting reactant - Fe; m(FeCl3) = 14.60 g